# Xinlu Zhang1, Chenxin Tian2, Xianjun Yang1, Lichang Chen3, Zekun Li1, Linda Ruth Petzold 1

### v3

@@ 5.1 Main Results
Benchmark Evaluation Performance. Table 3 presents an extensive evaluation of AlpaCare on 5 medical benchmark. AlpaCare obtain best performance on average, highlighting its robust capability in the medical domain. Benchmarks evaluate a model’s intrinsic knowledge(Gao et al., 2021), which is mainly gained in LLM pretraining instead of instruction fine-tuning (Zhou et al., 2023). AlpaCare’s strong medical capability, combined with its superior ability to follow medical instructions, enables it to meet a wide range of medical application needs effectively.
@@ 5.2 Generalizability Evaluation
Training models with specific data may lead to catastrophic forgetting, limiting their generalizability (Kirkpatrick et al., 2017). Our approach, instruction tuning a model with a diverse, domain-specific dataset, aims to improve its generalizability simultaneously. We test this using AlpaCare in AlpaFarm (Dubois et al., 2023), MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022) and TruthfulQA (Lin et al., 2022). We compare AlpaCare with 4 reference LLMs in AlpaFarm and report the average score. The detailed score for each reference model is deferred to the Table 13 in the Appendix. We follow Chia et al. (2023) to holistically evaluate models’ general domain knowledge on MMLU (5-shot) and BBH (3-shot), receptively; and evaluate the model truthfulness on TruthfulQA (0-shot) with Gao et al. (2021). The results are shown in Table 4.
Table 4: Performance on general domain tasks. AlpacaFarm is a free-form instruction evaluation, MMLU and BBH are knowledge benchmarks and TruthfulQA is truthfulness task. ‘AVG’ denotes the average score across all tasks.
@@ 5.3 Ablation Study
AlpaCare achieves superior performance across various backbones. To explore the effect of different LLM backbones, we tune Alpaca-LLaMA2 and AlpaCare-LLaMA2 by training LLaMA2-7B on Alpaca data and MedInstruct-52k, respectively. Table 6 compares performance of Alpaca and AlpaCare based on different LLM backbone families.
@@
AlpaCare shows robust performance across different judges. Recent studies have highlighted potential biases in the LLM evaluator (Wang et al., 2023a). ChatGPT may give higher preference on outputs from ChatGPT and GPT-4, which both trained by OpenAI. To robustly evaluate our method, we introduce an alternative judge, Claude-2 (Anthropic, 2023) from Anthropic, to mitigate the potential biases of relying on a single family of judges. The results are shown in Table 7.
Upon evaluation by Claude-2, it is observed that AlpaCare consistently outperforms its IFT baselines with a large margin. This aligns with findings from assessments using GPT-3.5-turbo as the judge. Such consistency underscores the superior medical proficiency of our approach.
@@ 6 Human Study
We further conduct human studies to label question-and-answer pairs in medical free-form instruction evaluation. Three annotators with MD degree in progress are involved in the study to perform pairwise comparisons for each question and answer pair. Specifically, we randomly select 50 prompts from each test set, totaling 100 prompts. These prompts, along with the responses generated by both AlpaCare-13B and PMC-13B, the best baseline in the 13B models, are presented to the annotators for evaluation. The evaluation is based on two criteria: correctness and helpfulness. Correctness evaluates whether the response provides accurate medical knowledge to address question posed, while helpfulness measures the model’s ability to assist users concisely and efficiently, considering the user intent. In practical terms, an answer can be correct but not necessarily helpful if it is too verbose and lacks guidance. To determine the final result for each criterion of each evaluation instance, we employ a majority vote method. If at least two of the annotators share the same opinion, their preference is considered the final answer; otherwise, we consider the outputs of the two models to be tied. The results are shown in Figure 3.
Consistent with previous results, the AlpaCare-13B obtains better performance than PMC-13B in the human evaluation. The output generated by AlpaCare-13B is more accurate than that from PMC-13B, with 54% answers preferred by expert annotators. This result shows superior medical capacity of AlpaCare. Beyond correctness, AlpaCare generates answers that are significantly more helpful compared to PMC-13B. 69% answers generated by AlpaCare are preferred by domain experts, highly exceeding 11% preference rate for PMC-13B outputs, demonstrating better practical usability. The higher relative gain in helpfulness of AlpaCare comparing to correctness is expected because the goal of IFT is to obtain the instruction-following ability of LLMs to align the model with various user needs, rather than gaining new knowledge (Zhou et al., 2023).

### v4

@@ 5.1 Main Results
Benchmark Evaluation Performance. Table 3 presents an extensive evaluation of AlpaCare on 5 medical benchmarks. AlpaCare obtain the best performance on average, highlighting its robust capability in the medical domain. Benchmarks evaluate a model’s intrinsic knowledge(Gao et al., 2021), which is mainly gained in LLM pretraining instead of instruction fine-tuning (Zhou et al., 2023). AlpaCare’s strong medical capability, combined with its superior ability to follow medical instructions, enables it to meet a wide range of medical application needs effectively.
@@ 5.2 Generalizability Evaluation
Training models with specific data may lead to catastrophic forgetting, limiting their generalizability (Kirkpatrick et al., 2017). Our approach, instruction tuning a model with a diverse, domain-specific dataset, aims to improve its generalizability simultaneously. We test this using AlpaCare in AlpaFarm (Dubois et al., 2023), MMLU (Hendrycks et al., 2021), BBH (Suzgun et al., 2022) and TruthfulQA (Lin et al., 2022). We compare AlpaCare with 4 reference LLMs in AlpaFarm and report the average score. The detailed score for each reference model is deferred to Table 13 in the Appendix. We follow Chia et al. (2023) to holistically evaluate models’ general domain knowledge on MMLU (5-shot) and BBH (3-shot), receptively; and evaluate the model truthfulness on TruthfulQA (0-shot) with Gao et al. (2021). The results are shown in Table 4.
Table 4: Performance on general domain tasks. AlpacaFarm is a free-form instruction evaluation, MMLU and BBH are knowledge benchmarks and TruthfulQA is a truthfulness task. ‘AVG’ denotes the average score across all tasks.
@@ 5.3 Ablation Study
AlpaCare achieves superior performance across various backbones. To explore the effect of different LLM backbones, we tune Alpaca-LLaMA2 and AlpaCare-LLaMA2 by training LLaMA2-7B on Alpaca data and MedInstruct-52k, respectively. Table 6 compares the performance of Alpaca and AlpaCare based on different LLM backbone families.
@@
AlpaCare shows robust performance across different judges. Recent studies have highlighted potential biases in the LLM evaluator (Wang et al., 2023a). ChatGPT may give a higher preference for outputs from ChatGPT and GPT-4, which are both trained by OpenAI. To robustly evaluate our method, we introduce an alternative judge, Claude-2 (Anthropic, 2023) from Anthropic, to mitigate the potential biases of relying on a single family of judges. The results are shown in Table 7.
Upon evaluation by Claude-2, it is observed that AlpaCare consistently outperforms its IFT baselines by a large margin. This aligns with findings from assessments using GPT-3.5-turbo as the judge. Such consistency underscores the superior medical proficiency of our approach.
@@ 6 Human Study
We further conduct human studies to label question-and-answer pairs in medical free-form instruction evaluation. Three annotators with MD degrees in progress are involved in the study to perform pairwise comparisons for each question and answer pair. Specifically, we randomly select 50 prompts from each test set, totaling 100 prompts. These prompts, along with the responses generated by both AlpaCare-13B and PMC-13B, the best baseline in the 13B models, are presented to the annotators for evaluation. The evaluation is based on two criteria: correctness and helpfulness. Correctness evaluates whether the response provides accurate medical knowledge to address the question posed, while helpfulness measures the model’s ability to assist users concisely and efficiently, considering the user intent. In practical terms, an answer can be correct but not necessarily helpful if it is too verbose and lacks guidance. To determine the final result for each criterion of each evaluation instance, we employ a majority vote method. If at least two of the annotators share the same opinion, their preference is considered the final answer; otherwise, we consider the outputs of the two models to be tied. The results are shown in Figure 3.
Consistent with previous results, the AlpaCare-13B obtains better performance than PMC-13B in the human evaluation. The output generated by AlpaCare-13B is more accurate than that from PMC-13B, with 54% answers preferred by expert annotators. This result shows a superior medical capacity of AlpaCare. Beyond correctness, AlpaCare generates answers that are significantly more helpful compared to PMC-13B. 69% answers generated by AlpaCare are preferred by domain experts, highly exceeding the 11% preference rate for PMC-13B outputs, demonstrating better practical usability. The higher relative gain in helpfulness of AlpaCare compared to correctness is expected because the goal of IFT is to obtain the instruction-following ability of LLMs to align the model with various user needs, rather than gaining new knowledge (Zhou et al., 2023).